### v1

# SIM-Shapley: A Stable and Computationally Efficient Approach to Shapley Value Approximation

Abstract
Explainable artificial intelligence (XAI) is essential for trustworthy machine learning (ML), particularly in high-stakes domains such as healthcare and finance. Shapley value (SV) methods provide a principled framework for feature attribution in complex models but incur high computational costs, limiting their scalability in high-dimensional settings. We propose Stochastic Iterative Momentum for Shapley Value Approximation (SIM-Shapley), a stable and efficient SV approximation method inspired by stochastic optimization. We analyze variance theoretically, prove linear $Q$-convergence, and demonstrate improved empirical stability and low bias in practice on real-world datasets. In our numerical experiments, SIM-Shapley reduces computation time by up to 85% relative to state-of-the-art baselines while maintaining comparable feature attribution quality. Beyond feature attribution, our stochastic mini-batch iterative framework extends naturally to a broader class of sample average approximation problems, offering a new avenue for improving computational efficiency with stability guarantees. Code is publicly available at https://github.com/nliulab/SIM-Shapley.
1 Introduction
As machine learning (ML) and artificial intelligence (AI) models grow increasingly complex, explainable AI (XAI) has emerged as a critical area for understanding model behavior and building trust in predictions doshivelez2017rigorousscienceinterpretablemachine ; covert2020understandingglobalfeaturecontributions . Quantifying feature importance is therefore essential in scientific and engineering domains where explainability is non-negotiable. One such domain is healthcare, where understanding the influence of individual predictors can inform targeted interventions and resource allocation ning2024variable .
The Shapley Value (SV) P-295 , originating in cooperative game theory, has become a foundation for XAI because it attributes a feature’s contribution to predictive performance while satisfying key fairness axioms lundberg2017unifiedapproachinterpretingmodel ; rodemann2024explainingbayesianoptimizationshapley ; 7546525 . SV-based explanations also support contrastive explanations, identify corrupted or influential features, and often align well with human intuition covert2020understandingglobalfeaturecontributions ; molnar2025 .
However, they incur exponential computational cost. Let $D=\{1,\dots,d\}$ be the index set of features and $S\subseteq D$ be a subset. A cooperative game is a set function $v:2^{d}\to\mathbb{R}$ that assigns a score to every subset $S$. To obtain the SV of feature $i\in D$, one must average its marginal contribution $v(S\cup\{i\})-v(S)$ over all $S$ that exclude $i$. Because there are $2^{d-1}$ such subsets for each feature, the total number of evaluations grow as $\mathcal{O}(d2^{d})$, which is impractical for large $d$.
To overcome this burden, we reformulate the problem as a stochastic optimization problem AguidetoSAA ; hannah2015stochastic . Our method, SIM-Shapley, iteratively refine estimates via mini-batch sampling and momentum updates polyakAverage ; RuppertAverage ; kingma2017adammethodstochasticoptimization ; tieleman2012rmsprop . An $\ell_{2}$ penalty ridge further reduces variance and enhances stability, yielding a flexible iterative framework broadly applicable to sample-average-approximation (SAA) problems. Our main contributions include:
1. A new SV estimator with provable linear $Q$-convergence, providing faster convergence than state-of-the-art methods while retaining consistency.
2. Stability mechanisms: $\ell_{2}$ regulation, detection of negative-sampling events, and correction of initialization bias.
3. A general, scalable framework compatible with diverse cooperative game formulations, naturally integrating techniques such as paired sampling covert2021improvingkernelshappracticalshapley , and extensible to a wider class of SAA objectives.
2 Related Work
2.1 Shapley Value as Weighted Least Squares
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ denote $d$ features for $n$ samples, and let $D=\{1,2,\dots,d\}$ represent their index set. Given a predictive model $f$ for the response $Y$, the performance of any subset $S\subseteq D$ can be expressed through a cooperative game $v(S)$. A restricted model $f_{S}=f(x_{S},X_{S^{c}})$ uses only features in $S$, with the complement $S^{c}=D\setminus S$ drawn from the conditional distribution $p(X_{S^{c}}\mid X_{S}=x_{S})$, yielding local explanations.
Several choices for $v(S)$ exist. SHAP lundberg2017unifiedapproachinterpretingmodel defines $v(S)$ as the expectation of the prediction from $f_{S}$, whereas Shapley Effects sveffects use output variance. In this work, we adopt the prediction-loss game:
$v(S)=-\textit{l}(\mathbb{E}[f_{S}\mid x_{S}],y)$ (1)
following SAGE covert2020understandingglobalfeaturecontributions . Besides local explanations, global explanations arise by treating an input-label pair $U=(X,Y)$ as exogenous covert2021improvingkernelshappracticalshapley and defining a stochastic cooperative game as:
$\mathbf{V}(S)=\mathbb{E}_{U}[v(S,U)],$ (2)
which captures the expected contribution of a feature subset across the entire data distribution covert2020understandingglobalfeaturecontributions . This formulation provides a unified framework that encompasses both local and global Shapley value–based methods.
The Shapley value P-295 (SV) $\phi_{i}$ of feature $i\in D$ is defined as the average marginal contribution of feature $i$ across all subsets $S$ that exclude $i$:
$\phi_{i}(v)=\frac{1}{d}\sum_{S\subseteq D\setminus\{i\}}\binom{d-1}{|S|}^{-1}\left(v(S\cup\{i\})-v(S)\right).$ (3)
Lundberg et al. lundberg2017unifiedapproachinterpretingmodel recast Shapley estimation as kernel-weighted linear regression, later refined by Covert et al.  covert2021improvingkernelshappracticalshapley . Approximating $v(S)$ with the additive form $u(S)=\beta_{0}+\sum_{i\in S}\beta_{i}$ (where $\beta_{0}=v(\emptyset),\sum_{i\in S}\beta_{i}=v(S)-v(\emptyset)$) leads to the weighted least squares objective:
$\min_{\beta_{0},\dots,\beta_{d}}\sum_{S\subseteq D}\mu_{\mathrm{Sh}}(S)\left(u(S)-v(S)\right)^{2},\qquad\mu_{\mathrm{Sh}}(S)=\frac{d-1}{\binom{d}{|S|}\,|S|\,(d-|S|)}$ (4)
where $\mu_{\text{Sh}}(S)$ is the Shapley kernel that yields the optimal solution equal to SV Charnes1988 ; lundberg2017unifiedapproachinterpretingmodel .
Let $\bm{\beta}=(\beta_{1},\dots,\beta_{d})$ denote the optimal coefficients (i.e., the minimizer) for Eq. (4); these are the estimated SV. To represent subsets more conveniently, map each subset $S\subseteq D$ to a binary vector $z\in\{0,1\}^{d}$, with $z_{i}=1\ \Leftrightarrow i\in S$. With a slight abuse of notation, write $v(z)=v(S)$ and $\mu_{\text{Sh}}(z)=\mu_{\text{Sh}}(S)$ for the subset $S=\{i:z_{i}=1\}$. We then define a sampling distribution $p(z)\propto\mu_{\text{Sh}}(z)$ over binary vectors $z$ satisfying $0<\mathbf{1}^{\top}z<d$; otherwise, we set $p(z)=0$.
Under this formulation, estimating SV reduces to solving the following constrained optimization problem covert2021improvingkernelshappracticalshapley :
$\min_{\beta_{1},\dots,\beta_{d}}\sum_{z}p(z)\left(v(\mathbf{0})+z^{\top}\bm{\beta}-v(z)\right)^{2}\quad\text{s.t.}\quad\mathbf{1}^{\top}\bm{\beta}=v(\mathbf{1})-v(\mathbf{0}).$ (5)
Since the squared term vanishes for $z=\mathbf{0}$ and $z=\mathbf{1}$, these configurations can be excluded from sampling.
2.2 Monte Carlo Estimation and Practical Challenges
Evaluating Eq. (5) on all $2^{d}$ subsets is computationally prohibitive. Instead, we draw $m$ i.i.d. samples $z_{i}\sim p(z)$ and their values $v(z_{i})$, yielding the Monte-Carlo surrogate:
$\min_{\beta_{1},\dots,\beta_{d}}\frac{1}{m}\sum_{i=1}^{m}\left(v(\mathbf{0})+z_{i}^{T}\bm{\beta}-v(z_{i})\right)^{2}\quad\text{s.t.}\quad\mathbf{1}^{T}\bm{\beta}=v(\mathbf{1})-v(\mathbf{0}).$ (6)
Covert et al. covert2021improvingkernelshappracticalshapley derive a closed-form solution via the Karush–Kuhn–Tucker boyd2004convex (KKT) conditions and proved consistency as $m\to\infty$. In practice, however, only a finite number of $z_{i}$ are available, which may lead to a singular system in certain regions of the solution space and cause the optimization to collapse. Moreover, naively averaging batch estimates yields high variance and fluctuation due to insufficient coverage of the subset space, leading to numerical instability and optimization failure.
2.3 From Sample Average Approximation to Stochastic Iteration
The constrained problem in Eq. (5) and its sampled version in Eq. (6) follow the Sample Average Approximation (SAA) paradigm AguidetoSAA ; shapiro1996convergence : we replace the expectation in Eq. (5) by the empirical average in Eq. (6). Under standard convexity and regularity conditions, the SAA solution converges almost surely as $m\to\infty$ fu2006gradient ; AguidetoSAA . However, large $m$ can be impractical due to computational constraints, and not all samples are equally informative for improving the estimator.
To address this, we turn to Stochastic Approximation (SA) methods hannah2015stochastic ; shapiro1996simulation , which update the solution incrementally using only a mini-batch at each step. This avoids repeatedly solving the full SAA problem and can lead to faster convergence in practice. As pointed out by Royset and Szechtman RePEc:inm:oropre:v:61:y:2013:i:3:p:762-776 , SA can outperform SAA in convergence rate, regardless of whether the underlying solver is sublinear, linear, or superlinear.
Moreover, sampling mini-batches of $z_{i}$ from the $2^{d}$ feature space inflates variance slows convergence qian2020impactminibatchsizevariance ; balles2017couplingadaptivebatchsizes . Adding an $\ell_{2}$ penalty stabilizes updates by shrinking large fluctuations in the estimator, reducing variance without breaking convexity fan2021communicationefficientaccuratestatisticalestimation ; shamir2014communicationefficientdistributedoptimization .
In Section 3, we introduce SIM-Shapley which combines stochastic momentum, adaptive averaging, and $\ell_{2}$ regularization to accelerate convergence and enhance numerical stability in SV estimation.
3 Proposed Method: SIM-Shapley
3.1 Stochastic Iteration with Momentum
Following the stochastic optimization formulation discussed in Section 2.3, we reformulate the KernelSHAP approximation lundberg2017unifiedapproachinterpretingmodel in Eq. (6) as an iterative procedure to update estimates $\bm{\beta}^{(n)}$ using an exponential moving average (EMA) scheme:
$\min_{\beta^{(n+1)}_{1},\dots,\beta^{(n+1)}_{d}}$ $\frac{1}{m}\sum_{i=1}^{m}\left(v(\mathbf{0})+z_{i}^{T}\bm{\beta}^{(n+1)}-v(z_{i})\right)^{2}+\lambda\left\|\bm{\delta}^{(n+1)}\right\|_{2}^{2}$ (7a)
s.t. $\mathbf{1}^{T}\bm{\beta}^{(n+1)}=v(\mathbf{1})-v(\mathbf{0})$
$\bm{\beta}^{(n+1)}=t\bm{\beta}^{n}+(1-t)\bm{\delta}^{(n+1)}$ (7b)
where $10<t<$ 1 is a fixed momentum parameter. $\bm{\delta}^{(n)}$ represents the sampling information obtained at the $n$-th iteration, which can be interpreted as a form of momentum. The updated estimator $\bm{\beta}^{(n)}$ is constructed as a weighted combination of the previous estimate $\bm{\beta}^{(n-1)}$ and $\bm{\delta}^{(n)}$. In addition, $\bm{\beta}^{(n)}$ shall satisfy the constraint in Eq. (5) to preserve the additive property of SV, as reflected in the definition of $u(S)$ (Eq. (4)).
To control the variance of $\bm{\delta}^{(n)}$ under limited sampling (e.g., when $\{z_{1},\dots,z_{m}\}$ contains many duplicate subsets), we introduce an $\ell_{2}$ regularization penalty with coefficient $\lambda>0$. This penalty ensures strict convexity of the objective function at each iteration and guarantees the existence and uniqueness of a closed-form solution. Intuitively, the penalty stabilizes the updates by discouraging large deviations of $\bm{\delta}^{(n)}$ from $\bm{\beta}^{(n-1)}$.
For notational simplicity, we do not distinguish between $v(S)$ in Eq. (1) and $\mathbf{V}(S)$ in Eq. (2), as replacing $v(S)$ with $\mathbf{V}(S)$ allows $\bm{\beta}^{(n+1)}$ in SIM-Shapley to be interpreted as an estimator of the global SV explanation. To approximate $\mathbf{V}(S)$, we construct a Global Reference Set by sampling from the original dataset, thereby approximating the distribution $p(U)$ and enabling the evaluation of $\mathbb{E}_{U}[v(S,U)]$ for each cooperative game.
As discussed in Section 2.1, to approximate the conditional distribution $p(X_{S^{c}}\mid X_{S}=x_{S})$, we handle the missing features $X_{S^{c}}$ by marginalizing them using their joint marginal distribution. Specifically, we build a Background Set from the data and use it to impute the removed variables. Since our focus is on estimating the SV rather than the underlying games themselves, this treatment does not affect the convergence or correctness of the iterative procedure.
Building on these ideas, we propose the SIM-Shapley algorithm, together with its local explanation version described in Algorithm 1. The global version is provided in the Appendix B.
Next we apply the Karush–Kuhn–Tucker (KKT) conditions boyd2004convex to Eq. (7), leveraging the convexity of the objective function and the linearity of the constraint. Given a sampled batch $(z_{1},\dots,z_{m})$, the associated Lagrangian function, with Lagrange multiplier $\nu\in\mathbb{R}$, is defined as follows:
$\mathcal{L}(\bm{\delta},\nu)=\frac{1}{m}\sum_{i=1}^{m}\left(v(\mathbf{0})+z_{i}^{\top}\bm{\beta}^{(n+1)}-v(z_{i})\right)^{2}+\lambda\|\bm{\delta}^{(n+1)}\|_{2}^{2}+\nu\left(\mathbf{1}^{\top}\bm{\delta}^{(n+1)}-c\right).$ (8)
where we further introduce the following shorthand notations:
$A$ $=\frac{1}{m}\sum_{i=1}^{m}z_{i}z_{i}^{\top},$ $\quad\bar{A}$ $=A+\lambda I,$ $\quad\bar{b}$ $=\frac{1}{m}\sum_{i=1}^{m}z_{i}\left(v(z_{i})-v(\mathbf{0})\right),$ $\quad c$ $=v(\mathbf{1})-v(\mathbf{0}).$
Solving the KKT conditions yields the closed-form update for $\bm{\delta}^{(n+1)}$:
$\bm{\delta}^{(n+1)}=\frac{1}{1-t}\bar{A}^{-1}\left[\left(\bar{b}-tA\bm{\beta}^{(n)}\right)+\mathbf{1}\cdot\frac{c-t\mathbf{1}^{\top}\bm{\beta}^{(n)}-\mathbf{1}^{\top}\bar{A}^{-1}\left(\bar{b}-tA\bm{\beta}^{(n)}\right)}{\mathbf{1}^{\top}\bar{A}^{-1}\mathbf{1}}\right]$ (9)
3.2 Variance and Convergence
We further analyze the variance and convergence behavior of the SIM-Shapley estimator $\bm{\beta}^{(n)}$ with respect to the ideal Shapley value vector $\bm{\beta}$. Throughout this section, we treat the KernelSHAP estimator in Eq. (6) as the ground truth, consistent with prior work demonstrating that KernelSHAP and its variants are either unbiased or nearly unbiased in practice covert2021improvingkernelshappracticalshapley ; covert2020understandingglobalfeaturecontributions ; chen2022algorithmsestimateshapleyvalue ; lundberg2017unifiedapproachinterpretingmodel . Since SIM-Shapley is designed to approximate KernelSHAP through a stochastic iterative scheme with regularization, any potential bias is empirically negligible, as confirmed by our experiments.
The total mean squared error of $\bm{\beta}^{(n)}$ admits the standard decomposition:
$\mathrm{MSE}\bigl{(}\bm{\beta}^{(n)}\bigr{)}\;:=\;\mathbb{E}\bigl{[}\|\bm{\beta}^{(n)}-\bm{\beta}\|_{2}^{2}\bigr{]}\;=\;\underbrace{\mathbb{E}\bigl{[}\|\bm{\beta}^{(n)}-\mathbb{E}[\bm{\beta}^{(n)}]\|_{2}^{2}\bigr{]}}_{\text{Variance}}\;+\;\underbrace{\|\mathbb{E}[\bm{\beta}^{(n)}]-\bm{\beta}\|_{2}^{2}}_{\text{Bias}^{2}}\,.$ (10)
In the following, we focus on the variance and convergence properties of SIM-Shapley across iterations, as they dominate the practical performance in typical settings where bias is negligible as expected, as later shown in Section 5.
Theorem 1 (Variance Contraction).
Let $\hat{\bm{\beta}}$ denote the estimator obtained from KernelSHAP in Eq. (6) with $m$ samples of $z$, and let $\bm{\beta}^{(n)}$ be the SIM-Shapley iterates with fixed momentum $t\in(0,1)$. As $m\to\infty$, for fixed momentum $t\in(0,1)$ and any feature index $i$ and any iteration $n\geq 1$,
$\operatorname{Var}(\beta^{(n)}_{i})\leq(1-t)^{2}\operatorname{Var}(\hat{\beta}_{i}).$
A detailed proof of Theorem 1 is given in Appendix A.1. In essence, each EMA update in SIM-Shapley reduces the variance of feature $i$ by at least a factor of $(1-t)^{2}$ relative to the vanilla SAA estimator $\hat{\bm{\beta}}_{i}$. Because the bias remains comparable to KernelSHAP, this geometric variance contraction provides a principled way to decide when to stop iterating.
Concretely, we monitor the largest standard error across all coordinates and compare it to the span of the current Shapley estimates. Denoting by $\sigma^{(n)}_{i}\;=\;\sqrt{\frac{\operatorname{Var}(\beta^{(n)}_{i})}{n}},$ the standard error of feature $i$ at iteration $n$, we declare convergence once
$\max_{i}\sigma^{(n)}_{i}\;<\;\epsilon\;\bigl{(}\max_{i}\beta^{(n)}_{i}-\min_{i}\beta^{(n)}_{i}\bigr{)},$ (11)
where $\epsilon$ is a small tolerance (e.g. 0.025). This criterion ensures that the residual uncertainty in any coordinate is a negligible fraction of the overall range of SV. To track $\sigma_{i}^{(n)}$ online with negligible extra memory, we employ Welford’s algorithm Welford01081962 within Algorithm 1.
Due to the strict convexity of the underlying quadratic objective and the affine constraint structure, SIM-Shapley admits guaranteed convergence. We now characterize its rate of convergence in terms of the $Q$-linear convergence criterion.
Theorem 2.
As $m\to\infty$, SIM-Shapley exhibits a linear Q-convergence rate, i.e.,
$\|\bm{\beta}^{(n)}-\bm{\beta}^{*}\|=\mathcal{O}(\rho(G)^{n}),$
where $\bm{\beta}^{*}$ is the optimal point of SIM-Shapley, $\rho(G)=t\cdot\frac{\lambda}{\alpha+\lambda}$, and $\alpha>0$ denotes the minimal eigenvalue of $A$.
Since $10<t<$ 1 and $10<\tfrac{\lambda}{\alpha+\lambda}<$ 1, $\rho=t\,\frac{\lambda}{\alpha+\lambda}$ satisfies $\rho<1$, guaranteeing linear convergence of SIM-Shapley.
Theorem 2 reveals a fundamental trade-off between convergence speed and bias, governed by the choice of $t$ and $\lambda$. The detailed proof is in Appendix A.2. In practice, our objective is accurate SV approximation rather than exact fixed-point attainment, so we continue to use the variance-based stopping rule of Eq. (11) for early termination.
4 Stability Improvement
In this section, we introduce two techniques to enhance the convergence stability of SIM-Shapley. By combining these methods, we propose Stable-SIM-Shapley (see Appendix C for complete pseudocode), an improved variant with more robust convergence behavior.
4.1 Preventing Negative Sampling
As shown in Section 3.2, the convergence of $\bm{\beta}^{(n+1)}$ is driven by the variance (covariance) of the update $\bm{\delta}^{(n+1)}$. In practice, the convergence of SIM-Shapley can stall or even reverse when a mini-batch contains many low-quality or duplicate coalitions—what we call negative sampling. Such batches produce high-variance updates $\bm{\delta}^{(n+1)}$, slowing overall convergence.
To detect and avoid these harmful updates, we compare the update variance to the current iterate variance. Specifically, we compute
$r=\frac{\left\|\mathrm{Var}(\bm{\delta}^{(n+1)})\right\|-\left\|\mathrm{Var}(\bm{\beta}^{(n)})\right\|}{\left\|\mathrm{Var}(\bm{\beta}^{(n)})\right\|}\leq\xi,$ (12)
and if
$r\;>\;\xi,\quad 0<\xi<1,$
we reject the batch and resample. In our experiments, $\xi\in\{0.30,0.35\}$ works well.
Discarding a single batch incurs only the marginal cost of one additional mini-batch draw, since we already track variances for convergence detection. This simple check dramatically reduces the frequency of high-variance updates, yielding smoother, faster convergence in practice.
4.2 Correcting Initialization Bias
The update rule in our algorithm follows an EMA scheme. However, such schemes are sensitive to initialization and typically suffer from initial bias. Specifically, during early iterations, $\bm{\beta}^{(n)}$ is overly influenced by $\bm{\beta}^{(0)}$, which holds a disproportionately large weight.
To address this, we initialize $\bm{\beta}^{(0)}=0$ and apply a bias correction technique similar to that used in the Adam optimizer kingma2017adammethodstochasticoptimization . This leads to a dynamically adjusted update rule, where the weight factor $t$ varies over iterations. The closed-form solution in Eq. (9) can be adjusted accordingly due to linearity (see Appendix C). The bias-corrected update rule is given by:
$\bm{\beta}^{(n+1)}=\frac{t\bm{\beta}^{(n)}}{1-t^{n+1}}+\frac{(1-t)\bm{\delta}^{(n+1)}}{1-t^{n+1}}.$ (13)
Algorithm 1 SIM-Shapley (Local Explanation)
0:  Instance $(x,y)$, background set $(X_{\mathrm{BK}},Y_{\mathrm{BK}})$, subset sample size $m$, number of iterations $T$
0:  Final estimate $\bm{\beta}^{(n)}$
1:  Initialize $\bm{\beta}^{(0)}=\mathbf{0}$; $Mean_{0}=\mathbf{0}$; $S_{0}=\mathbf{0}$; precompute $v(\mathbf{1})$ and $v(\mathbf{0})$ for the target instance
2:  for $n=1$ to $T$ do
3:     Independently sample $\{z_{j}\}_{j=1}^{m}\sim p(z)$
4:     For each $z_{j}$, compute $v(z_{j})$ by marginalizing missing features using the background set
5:     Solve Eq. (7) to obtain $\bm{\delta}^{(n)}$ and update $\bm{\beta}^{(n)}$
6:     Update running variance $Var_{n}$ of $\bm{\beta}^{(n)}$ using Welford’s algorithm:
7:      $\Delta\leftarrow\bm{\beta}^{(n)}-Mean_{n-1}$;  $Mean_{n}\leftarrow Mean_{n-1}+\Delta/n$
8:      $S_{n}\leftarrow S_{n-1}+\Delta\cdot(\bm{\beta}^{(n)}-Mean_{n})$;  $Var_{n}\leftarrow S_{n}/(n-1)$
9:     if convergence criterion satisfied then
10:        break
11:     end if
12:  end for
13:  return  $\bm{\beta}^{(n)}$
5 Experiments
5.1 Benchmark Results on Standard ML Tasks
We first validated our method on four standard machine learning datasets [^1] All experiments were conducted on a 12th Gen Intel Core i5-12600KF CPU and an NVIDIA RTX 2060 GPU. Each experiment was repeated 100 times.: the Bike Sharing Demand dataset bike , the South German Credit dataset south_german_credit_522 , the Portuguese Bank Marketing dataset bank_marketing_222 , and the NYC Airbnb Listings dataset gomonov2019nycairbnb . We evaluated performance using three predictive models: XGBoost xgboost10.1145/2939672.2939785 , CatBoost cat10.5555/3327757.3327770 , and a multi-layer perceptron (MLP) implemented in PyTorch ansel2024pytorch2 , as summarized in Tables 1 and 2. We benchmarked our method against SAGE covert2020understandingglobalfeaturecontributions for global explanations and KernelSHAP lundberg2017unifiedapproachinterpretingmodel for local explanations, both of which are approximation methods for the objective in Eq. (4).
Table 1: Global explainability benchmarks. Average running time (seconds) and consistency (Pearson correlation with baseline) of SIM-Shapley and Stable-SIM-Shapley across four datasets.
Data Model Avg. Running Time (s) Avg. Consistency
SIM-Shapley Stable-SIM-Shapley SAGE SIM-Shapley Stable-SIM-Shapley
Bike XGBoost 7.95 7.56 21.66 0.9958 0.9966
Credit CatBoost 23.88 17.59 61.39 0.9654 0.9675
Bank CatBoost 46.39 37.16 171.39 0.9529 0.9518
Airbnb MLP 7.05 4.87 32.80 0.9588 0.9659
Table 2: Local explainability benchmarks. Average running time (seconds) and consistency (Pearson correlation with baseline) of SIM-Shapley and Stable-SIM-Shapley across four datasets.
Data Model Avg. Running Time (s) Avg. Consistency
SIM-Shapley Stable-SIM-Shapley KernelSHAP SIM-Shapley Stable-SIM-Shapley
Bike XGBoost 8.48 8.59 12.62 0.9594 0.9594
Credit CatBoost 0.52 0.45 3.11 0.9831 0.9829
Bank CatBoost 1.73 2.30 7.49 0.9816 0.9816
Airbnb MLP 1.76 1.62 3.20 0.9111 0.9121
For fair comparison, all methods approximate the conditional distribution $p(X_{\bar{S}}\mid X_{S}=x_{S})$ using marginal distributions estimated from the background dataset (see Section 3.1). Mean squared error (MSE) is used for regression tasks (i.e., Airbnb and Bike) and cross-entropy (CE) is used for classifications (i.e., Bank and Credit). For each experiment on the same dataset and explanation, we report the average wall-clock time under a unified convergence threshold. Detailed settings are provided in Appendix D.2.
Our proposed method yields estimates that are consistent with two state-of-the-art baselines while significantly improving computational efficiency. Consistency is assessed using Pearson’s correlation coefficient, which is appropriate given the approximately homogeneous distribution of Shapley estimates. A detailed description and corresponding figures are provided in Appendix D.1 to further illustrate this.
We evaluate the convergence behavior of our methods against baseline approaches. As shown in Figure 1, in each single experiments, Stable-SIM-Shapley converges more smoothly than SIM-Shapley due to its enhanced stability mechanisms. Both variants effectively correct initial bias and achieve faster convergence without compromising consistency (Figure 2, Tables 1 and 2). Complete results are provided in Appendix D.2.
(a) Credit - Global Explanation
(b) Credit - Local Explanation
(c) Airbnb - Global Explanation
(d) Airbnb - Local Explanation
Figure 1: Comparison of convergence time (in seconds) on the Credit and Airbnb datasets. All methods incur a nonzero initialization time due to the computational cost of the first iteration. KernelSHAP may lack a valid convergence ratio during early iterations.
(a) Airbnb - Global - Consistency
(b) Airbnb - Global - Running Time
(c) Bank - Local - Consistency
(d) Bank - Local - Running Time
Figure 2: Boxplots of consistency (Pearson correlation) and running time (50 trials) for both global (Airbnb) and local (Bank) explainability settings. SIM-Shapley and Stable-SIM-Shapley are compared against baseline methods on both datasets.
5.2 Real-World Interpretability Assessment
To validate our methods in a realistic clinical setting, we apply SIM-Shapley and Stable-SIM-Shapley to the MIMIC-IV-ED dataset (johnson2023mimic, ), focusing on in-hospital mortality prediction (see Appendix D.3.1 for dataset details). In addition to quantitative results (Section 5.1), we incorporate expert review from a board-certified emergency physician, who confirms that the top-ranked features–such as vital signs, age, and comorbidity burden–are aligned with established medical knowledge. This supports the interpretability and practical utility of our methods. Full commentary appears in Appendix D.3.2.
Table 3: Top-10 features identified by each SV method on the MIMIC cohort. A check mark (✓) indicates inclusion in the method’s top-10 ranking. Results are shown for both global and local explanations.
Feature Global Explanation Local Explanation
SAGE SIM-Shapley Stable-SIM KernelSHAP SIM-Shapley Stable-SIM
Abdominal Pain
AIDS
Age
Congestive heart failure
Diabetes with complications
Diastolic blood pressure
Drug abuse
Fever/chills
Fluid and electroyte disorders
Gender
Heart rate at triage
Hemiplegia
Hypertension–complicated
Hypertension–uncomplicated
Local tumor, leukemia and lymphoma
Metastatic solid tumor
Myocardial infarction
Number of ED admission within the past year
Number of hospitalizations within the past year
Obesity
Peripheral oxygen saturation
Peripheral vascular disease
Psychoses
Respiration rate at triage
Stroke
Systolic blood pressure
Syncope
Temperature at triage
Weight Loss
6 Discussion
We have shown that reframing SV estimation as a stochastic optimization task—combining mini-batch sampling with momentum-driven updates and $\ell_{2}$ regularization—yields substantial gains in both speed and stability. Our SIM-Shapley algorithm converges linearly to the true solution (Theorem 2), with provable variance reduction at each step (Theorem 1), and registers negligible bias (Section 3.2). Empirically, on both standard benchmarks and the MIMIC-IV-ED clinical cohort, SIM-Shapley and its Stable variant reduce compute time by up to 85% while producing explanations that align with domain knowledge.
Crucially, our approach preserves the original weighted least-squares formulation, so all existing KernelSHAP enhancements—unbiased $A$ estimators, paired-sampling schemes, alternative cooperative-game definitions, and uncertainty quantification—can be dropped seamlessly into SIM-Shapley. The addition of an $\ell_{2}$ penalty strengthens convexity and invertibility, incurring no extra algorithmic complexity.
More broadly, our stochastic-iterative framework generalizes to any sampling-dependent quadratic minimization, offering a unifying foundation for accelerating Monte Carlo–based approximation tasks. Future work could explore adaptive momentum scaling, richer variance-reduction strategies, and extensions to other cooperative-game objectives beyond SV.
Like many existing SV approximation techniques, our method also adopts the marginal distribution as a proxy for the true conditional distribution, which may introduce approximation error in practice. Future work could explore more accurate conditional estimators—such as those based on deep generative models or copula methods—to improve attribution reliability. Additionally, it would be interesting to investigate hybrid frameworks that directly incorporate conditional sampling into the SV estimation process while preserving computational efficiency.
Appendix A Complete Proofs
Remark 1.
As $m\to\infty$, the empirical matrix $A$ converges in probability to a positive definite population matrix $\Sigma:=\mathbb{E}[z_{i}z_{i}^{\top}]$, whose minimum eigenvalue has lower bound $\frac{1}{4}$.
Proof.
Recall that the matrix $A$ is defined as:
$A=\frac{1}{m}\sum_{i=1}^{m}z_{i}z_{i}^{\top},$
where each $z_{i}\in\{0,1\}^{d}$ is drawn independently and uniformly from $\mathbb{R}^{d}$.
Let $\Sigma:=\mathbb{E}[z_{i}z_{i}^{\top}]$ denote the population second-moment matrix. Since the $z_{i}$ are i.i.d. and bounded random vectors, applying the law of large numbers element-wise gives us:
[…]
Since $z_{i}$ is uniformly distributed over $\{0,1\}^{d}$, its coordinates are independent and identically distributed Bernoulli($\frac{1}{2}$). Then
for diagonal entries ($k=\ell$):
$\mathbb{E}[z_{ik}^{2}]=\mathbb{E}[z_{ik}]=\frac{1}{2};$
for off-diagonal entries ($k\neq\ell$):
$\mathbb{E}[z_{ik}z_{i\ell}]=\mathbb{E}[z_{ik}]\mathbb{E}[z_{i\ell}]=\frac{1}{4}.$
Thus, the population matrix takes the explicit form:
$\Sigma=\mathbb{E}[z_{i}z_{i}^{\top}]=\frac{1}{4}\mathbf{1}_{d}\mathbf{1}_{d}^{\top}+\frac{1}{4}I_{d}.$
Multiplying a non-zero vector $v\in\mathbb{R}^{d}$ on the right hand side gives us:
$\mathbb{E}[z_{i}z_{i}^{T}]v=\left(\frac{1}{4}\mathbf{1}_{d}\mathbf{1}_{d}^{T}+\frac{1}{4}I_{d}\right)v=\frac{1}{4}\mathbf{1}_{d}(\mathbf{1}_{d}^{T}v)+\frac{1}{4}v$
As a result, its eigenvalues are:
  • $\lambda_{1}=\frac{1}{4}(d+1)$, corresponding to the direction $\mathbf{1}_{d}$,
  • $\lambda_{2}=\cdots=\lambda_{d}=\frac{1}{4}$, corresponding to any direction orthogonal to $\mathbf{1}_{d}$.
and $\alpha:=\text{min}(\lambda_{1},\lambda_{2})=\frac{1}{4}$. ∎
A.1 Poof of Theorem 1
Proof.
Recall the estimator of KernelSHAP in [16] is defined as $\bm{\beta}$ and the recursive form of $\bm{\beta}^{(n+1)}$ is as following:
$\bm{\beta}^{(n+1)}$ $=(1-t)\sum_{j=1}^{n+1}t^{n+1-j}\bm{\delta}^{(j)}.$
we see that each $\bm{\delta}^{(j)}$ depends on $\bm{\beta}^{(j-1)}$, making the iterations dependent through time.
We assume that the matrix $A=\frac{1}{m}\sum_{i=1}^{m}z_{i}z_{i}^{T}$ is positive definite with minimal eigenvalue $\alpha>0$, which holds asymptotically as $m\to\infty$ by Remark 1. This guarantees that the regularized matrix $\bar{A}=A+\lambda I$ is strictly positive definite, and $\bar{A}^{-1}$ is well-conditioned and bounded. Consequently, the stochastic updates remain well-behaved and the propagation of variance across iterations can be controlled.
The variance propagates recursively as:
$\operatorname{Var}(\bm{\beta}^{(n+1)})=t^{2}\operatorname{Var}(\bm{\beta}^{(n)})+(1-t)^{2}\operatorname{Var}(\bm{\delta}^{(n+1)})+2t(1-t)\operatorname{Cov}(\bm{\beta}^{(n)},\bm{\delta}^{(n+1)}).$
Assuming initialization $\bm{\beta}^{(0)}=\mathbf{0}$ and large $m$, the first step gives:
$\operatorname{Var}(\bm{\beta}^{(1)})=(1-t)^{2}\operatorname{Var}(\bm{\delta}^{(1)})\leq(1-t)^{2}\operatorname{Var}(\bm{\beta}).$
For subsequent steps, if $\operatorname{Var}(\bm{\delta}^{(j)})\leq\operatorname{Var}(\bm{\beta})$ (which holds under $\ell_{2}$ regularization), and the cross-covariance remains bounded, then:
$\operatorname{Var}(\bm{\beta}^{(2)})$ $=t^{2}(1-t)^{2}\operatorname{Var}(\bm{\delta}^{(1)})+(1-t)^{2}\operatorname{Var}(\bm{\delta}^{(2)})+2t(1-t)^{2}\operatorname{Cov}(\bm{\delta}^{(1)},\bm{\delta}^{(2)})$
$\leq(1-t)^{2}\operatorname{Var}(\bm{\delta}^{(1)})\leq(1-t)^{2}\operatorname{Var}(\bm{\beta}),$
$\vdots$
$\operatorname{Var}(\bm{\beta}^{(n+1)})$ $\leq(1-t)^{2}\max_{j\leq n+1}\operatorname{Var}(\bm{\delta}^{(j)})\leq(1-t)^{2}\operatorname{Var}(\bm{\beta}).$
We emphasize that $\bm{\delta}^{(n+1)}$ is computed conditional on $\bm{\beta}^{(n)}$, so independence does not hold. The covariance term $\operatorname{Cov}(\bm{\beta}^{(n)},\bm{\delta}^{(n+1)})$ cannot be assumed to vanish, but is expected to remain bounded under regularized updates. By the Cauchy-Schwarz inequality:
$|\operatorname{Cov}(\bm{\beta}^{(n)},\bm{\delta}^{(n+1)})|\leq\sqrt{\operatorname{Var}(\bm{\beta}^{(n)})\cdot\operatorname{Var}(\bm{\delta}^{(n+1)})},$
thus the recursion remains stable if both variances are controlled.
In conclusion, the variance of the SIM-Shapley estimator $\bm{\beta}^{(n+1)}$ is at most of the same order as that of the KernelSHAP estimator $\bm{\beta}$, and under ideal conditions may even be strictly smaller. This confirms that the variance reduction effect of the $\ell_{2}$ regularization is preserved across iterations. ∎
A.2 Poof of Theorem 2
Proof.
Under the assumption that $m\to\infty$ (Remark 1), we first prove the existence of optimal solution of SIM-Shapley, then analyze its quotient convergence rate,
A.2.1 Existence of a Fixed Point
Given a vector $\bm{\beta}$, define $\delta^{*}(\bm{\beta})$ as the unique solution to the population-level version of problem Eq. (7), where all empirical averages are replaced by their exact expectations. Define the mapping:
$T:\mathbb{R}^{d}\to\mathbb{R}^{d},\quad T(\bm{\beta})=t\,\bm{\beta}+(1-t)\,\delta^{*}(\bm{\beta}).$
A vector $\bm{\beta}^{*}$ is a fixed point if
$\bm{\beta}^{*}=T(\bm{\beta}^{*})=t\,\bm{\beta}^{*}+(1-t)\,\delta^{*}(\bm{\beta}^{*}).$ (14)
This implies directly that
$\delta^{*}(\bm{\beta}^{*})=\bm{\beta}^{*},$
i.e., the fixed point $\bm{\beta}^{*}$ satisfies the optimality condition of the constrained quadratic program.
Since the objective function in Eq. (7) is strictly convex and smooth, and the constraint is affine, the solution mapping $\delta^{*}(\bm{\beta})$ is continuous in $\bm{\beta}$, implying that $T$ is continuous. Moreover, due to the quadratic growth of the objective, there exists a compact, convex set $\mathcal{K}\subset\mathbb{R}^{d}$ (e.g., a sufficiently large closed ball) such that $T(\mathcal{K})\subset\mathcal{K}$.
Applying Brouwer’s fixed-point theorem to the continuous self-map $T:\mathcal{K}\to\mathcal{K}$, we conclude that a fixed point $\bm{\beta}^{*}\in\mathcal{K}$ exists.
A.2.2 Local Linearization and Convergence Rate via the Jacobian
To analyze the convergence behavior near the fixed point $\bm{\beta}^{*}$, we linearize the mapping $T$ in a neighborhood of $\bm{\beta}^{*}$.
Let the error at iteration $n$ be:
$e^{(n)}=\bm{\beta}^{(n)}-\bm{\beta}^{*}.$
Since $\bm{\beta}^{*}$ is a fixed point, $T(\bm{\beta}^{*})=\bm{\beta}^{*}$. Using the differentiability of $\delta^{*}(\bm{\beta})$, we apply a first-order Taylor expansion around $\bm{\beta}^{*}$:
$\delta^{*}(\bm{\beta})\approx\delta^{*}(\bm{\beta}^{*})+M(\bm{\beta}-\bm{\beta}^{*}),$
where
$M=\left.\frac{\partial\delta^{*}(\bm{\beta})}{\partial\bm{\beta}}\right|_{\bm{\beta}=\bm{\beta}^{*}}$
is the Jacobian evaluated at $\bm{\beta}^{*}$. Since $\delta^{*}(\bm{\beta}^{*})=\bm{\beta}^{*}$, this simplifies to:
$\delta^{*}(\bm{\beta})\approx\bm{\beta}^{*}+M(\bm{\beta}-\bm{\beta}^{*}).$
Substituting into the update rule:
$\bm{\beta}^{(n+1)}=t\,\bm{\beta}^{(n)}+(1-t)\,\delta^{*}(\bm{\beta}^{(n)}),$
we obtain:
$\bm{\beta}^{(n+1)}$ $\approx t\,\bm{\beta}^{(n)}+(1-t)\left[\bm{\beta}^{*}+M(\bm{\beta}^{(n)}-\bm{\beta}^{*})\right]$
$=t\,\bm{\beta}^{(n)}+(1-t)\bm{\beta}^{*}+(1-t)M(\bm{\beta}^{(n)}-\bm{\beta}^{*}).$
Subtracting $\bm{\beta}^{*}$ from both sides yields:
$e^{(n+1)}\approx\left[tI+(1-t)M\right]e^{(n)}.$
Define the linear update operator $G:=tI+(1-t)M$. Then, the convergence behavior is governed by the spectral radius:
$\rho(G)=\max\{|\lambda|:\lambda\in\sigma(G)\}.$
If $\rho(G)<1$, the error converges linearly:
$\|e^{(n)}\|=O(\rho(G)^{n}).$
In our setting, the matrix $A$ is given by:
$A=\frac{1}{m}\sum_{i=1}^{m}z_{i}z_{i}^{T},$
and the regularized version is $\bar{A}=A+\lambda I$. Through differentiation of the closed-form expression for $\delta^{*}(\bm{\beta})$ under the equality constraint, we obtain the Jacobian:
$M=-\frac{t}{1-t}\,\bar{A}^{-1}\left[A-\mathbf{1}\cdot\frac{\mathbf{1}^{T}(\bar{A}^{-1}A-I)}{\mathbf{1}^{T}\bar{A}^{-1}\mathbf{1}}\right].$
Thus, the linear operator $G$ becomes:
$G=tI+(1-t)M=t\left[I-\bar{A}^{-1}A+\bar{A}^{-1}\mathbf{1}\cdot\frac{\mathbf{1}^{T}(\bar{A}^{-1}A-I)}{\mathbf{1}^{T}\bar{A}^{-1}\mathbf{1}}\right].$
In the subspace orthogonal to the constraint direction, assume $x$ is an eigenvector of $\bar{A}^{-1}A$ with eigenvalue $\mu=\frac{\alpha_{i}}{\alpha_{i}+\lambda}<1$, where $\alpha_{i}$ is an eigenvalue of $A$. Then,
$Gx\approx t(1-\mu)x.$
Hence, the spectral radius of $G$ is:
$\rho(G)=\max_{i}|t(1-\mu_{i})|=t\cdot\frac{\lambda}{\alpha_{\min}^{+}+\lambda},$
where $\alpha_{\min}^{+}$ denotes the smallest nonzero eigenvalue of $A$, which has a lower bound (Remark 1). Since $10<t<$ 1 and $\frac{\lambda}{\alpha_{i}+\lambda}<1$, we conclude $\rho(G)<1$, and the iteration converges linearly:
$\|e^{(n)}\|=O(\rho(G)^{n}).$
Appendix B SIM-Shapley: Global Explanation Version
We exhibit SIM-Shapley for global explanation in Algorithm 2.
Algorithm 2 SIM-Shapley (Global Explanation)
0:  Global reference set $(X_{\mathrm{REF}},Y_{\mathrm{REF}})$ as $p(U)$, background set $(X_{\mathrm{BK}},Y_{\mathrm{BK}})$, subset sample size $m$, batch size $B$, number of iterations $T$
0:  Final estimate $\bm{\beta}^{(n)}$
1:  Initialize $\bm{\beta}^{(0)}=\mathbf{0}$; $Mean_{0}=\mathbf{0}$; $S_{0}=\mathbf{0}$; precompute $V(\mathbf{1})$ and $V(\mathbf{0})$ by averaging over $(X_{\mathrm{REF}},Y_{\mathrm{REF}})$
2:  for $n=1$ to $T$ do
3:     Independently sample $\{z_{j}\}_{j=1}^{m}\sim p(z)$
4:     Sample $\{(x_{j},y_{j})\}_{j=1}^{B}\sim p(U)$
5:     For each $z_{j}$, estimate $V(z_{j})$ by averaging predictions on $\{(x_{j},y_{j})\}_{j=1}^{B}$, marginalizing missing features using the background set
6:     Solve Eq. (7) to obtain $\bm{\delta}^{(n)}$ and update $\bm{\beta}^{(n)}$
7:     Update running variance $Var_{n}$ of $\bm{\beta}^{(n)}$ using Welford’s algorithm:
8:      $\Delta\leftarrow\bm{\beta}^{(n)}-Mean_{n-1}$;  $Mean_{n}\leftarrow Mean_{n-1}+\Delta/n$
9:      $S_{n}\leftarrow S_{n-1}+\Delta\cdot(\bm{\beta}^{(n)}-Mean_{n})$;  $Var_{n}\leftarrow S_{n}/(n-1)$
10:     if convergence criterion satisfied then
11:        break
12:     end if
13:  end for
14:  return  $\bm{\beta}^{(n)}$
Appendix C Stable SIM-Shapley
In this part, we introduce Algorithm 3 and 4 that is equipped with stability improving methods proposed in our paper. New iteration form and explicit solution of $\bm{\delta}$ are as following:
New iteration form with correcting initial deviation:
$\min_{\beta^{(n+1)}_{0},\dots,\beta^{(n+1)}_{d}}\frac{1}{m}\sum_{i=1}^{m}\left(v(\mathbf{0})+z_{i}^{T}\bm{\beta}^{(n+1)}-v(z_{i})\right)^{2}+\lambda\left\|\bm{\delta}^{(n+1)}\right\|_{2}^{2}$ (15)
[…]
New explicit solution with correcting initial deviation:
$c_{1}$ $=\frac{t}{1-t^{n+1}},\quad c_{2}=\frac{1-t}{1-t^{n+1}}$ (16)
$\bm{\delta}^{(n+1)}$ $=\frac{\bar{A}^{-1}[\bigl{(}\bar{b}-c_{1}A\bm{\beta}^{(n)}\bigr{)}+\mathbf{1}\frac{\bigl{(}v(\mathbf{1})-v(\mathbf{0})\bigr{)}-c_{1}\mathbf{1}^{T}\bm{\beta}^{(n)}-\mathbf{1}^{T}\bar{A}^{-1}\bigl{(}\bar{b}-c_{1}A\bm{\beta}^{(n)}\bigr{)}}{\mathbf{1}^{T}\bar{A}^{-1}\mathbf{1}}]}{c_{2}}.$
Algorithm 3 Stable SIM-Shapley (Local Explanation)
0:  Instance $(x,y)$, background set $(X_{\mathrm{BK}},Y_{\mathrm{BK}})$, subset sample size $m$, number of iterations $T$
0:  Final estimate $\bm{\beta}^{(n)}$
1:  Initialize $\bm{\beta}^{(0)}=\mathbf{0}$; $Mean_{0}=\mathbf{0}$; $S_{0}=\mathbf{0}$; precompute $v(\mathbf{1})$ and $v(\mathbf{0})$ for the target instance
2:  for $n=1$ to $T$ do
3:     Independently sample $\{z_{j}\}_{j=1}^{m}\sim p(z)$
4:     For each $z_{j}$, compute $v(z_{j})$ by marginalizing missing features using the background set
5:     Solve Eq. (15) to obtain $\bm{\delta}^{(n)}$ and update $\bm{\beta}^{(n)}$
6:     Update running variance $Var_{n}$ of $\bm{\beta}^{(n)}$ using Welford’s algorithm:
7:      $\Delta\leftarrow\bm{\beta}^{(n)}-Mean_{n-1}$;  $Mean_{n}\leftarrow Mean_{n-1}+\Delta/n$
8:      $S_{n}\leftarrow S_{n-1}+\Delta\cdot(\bm{\beta}^{(n)}-Mean_{n})$;  $Var_{n}\leftarrow S_{n}/(n-1)$
9:     if convergence criterion satisfied then
10:        break
11:     end if
12:     if negative sampling detected then
13:        Let $\bm{\delta}^{(n)}=\bm{\delta}^{(n-1)},\bm{\beta}^{(n)}=\bm{\beta}^{(n-1)}$
14:     end if
15:  end for
16:  return  $\bm{\beta}^{(n)}$
Algorithm 4 Stable SIM-Shapley (Global Explanation)
0:  Global reference set $(X_{\mathrm{REF}},Y_{\mathrm{REF}})$ as $p(U)$, background set $(X_{\mathrm{BK}},Y_{\mathrm{BK}})$, subset sample size $m$, batch size $B$, number of iterations $T$
0:  Final estimate $\bm{\beta}^{(n)}$
1:  Initialize $\bm{\beta}^{(0)}=\mathbf{0}$; $Mean_{0}=\mathbf{0}$; $S_{0}=\mathbf{0}$; precompute $V(\mathbf{1})$ and $V(\mathbf{0})$ by averaging over $(X_{\mathrm{REF}},Y_{\mathrm{REF}})$
2:  for $n=1$ to $T$ do
3:     Independently sample $\{z_{j}\}_{j=1}^{m}\sim p(z)$
4:     Sample $\{(x_{j},y_{j})\}_{j=1}^{B}\sim p(U)$
5:     For each $z_{j}$, estimate $V(z_{j})$ by averaging predictions on $\{(x_{j},y_{j})\}_{j=1}^{B}$, marginalizing missing features using the background set
6:     Solve Eq. (15) to obtain $\bm{\delta}^{(n)}$ and update $\bm{\beta}^{(n)}$
7:     Update running variance $Var_{n}$ of $\bm{\beta}^{(n)}$ using Welford’s algorithm:
8:      $\Delta\leftarrow\bm{\beta}^{(n)}-Mean_{n-1}$;  $Mean_{n}\leftarrow Mean_{n-1}+\Delta/n$
9:      $S_{n}\leftarrow S_{n-1}+\Delta\cdot(\bm{\beta}^{(n)}-Mean_{n})$;  $Var_{n}\leftarrow S_{n}/(n-1)$
10:     if convergence criterion satisfied then
11:        break
12:     end if
13:     if negative sampling detected then
14:        Let $\bm{\delta}^{(n)}=\bm{\delta}^{(n-1)},\bm{\beta}^{(n)}=\bm{\beta}^{(n-1)}$
15:     end if
16:  end for
17:  return  $\bm{\beta}^{(n)}$
Appendix D Additional Experimental Details
D.1 Pearson’s Correlation Coefficient as a Consistency Metric
In the main paper, we introduced several methods for approximating Shapley values. Under a common cooperative game formulation, these estimates should exhibit consistency, even though individual values may vary due to randomness in the sampling process. That is, while the exact numerical outputs may differ, the overall distribution of estimated Shapley values is expected to remain stable across methods.
To evaluate this consistency, prior work has employed regression analysis. As illustrated in Figure 1, the estimates from different methods tend to align closely along a straight line, indicating a strong linear relationship. This observation motivates the use of Pearson’s correlation coefficient as a quantitative measure of consistency. A coefficient near  11\,1\,1 suggests near-perfect linear agreement between two sets of Shapley estimates.
Moreover, when using the same cooperative game and an identical strategy for approximating the conditional distribution of missing features, the imposed equality constraint ensures that any bias remains bounded.
eFigure 1: Comparison of consistency across different Shapley value estimators. All axes represent estimated Shapley values. (a) Comparison between two SAGE estimates obtained via permutation and kernel regression. (b) Comparison between the SAGE estimate and the SIM-Shapley estimate. Experiments are conducted on the Bank dataset under identical settings. (c) Comparison between SIM-Shapley and its stable version. All SV estimates stem from Bike dataset with global explanation.
D.2 ML Tasks Details
We provide supplementary details for the experimental setup described in the main paper. Each dataset is split into training, validation, and test sets with a ratio of 7:2:1. Table 1 summarizes the dataset statistics and experimental configurations.
eTable 1: Dataset profile and experiment settings. $t$: the hyperparameter in update formula, which may be different in global and local explanation.
Dataset Total Size Feature Dimension Background Set Size Global Reference Set Size $\mathbf{z}$ Sample Size Batch Size $\mathbf{t}$ (Global/Local)
Bike 10886 12 512 1088 120 512 0.55 / 0.50
Credit 1000 20 256 100 200 512 0.55 / 0.55
Bank 45210 16 512 2000 160 512 0.55 / 0.50
Airbnb 48590 15 1024 4859 150 512 0.55 / 0.50
MIMIC-IV 414537 56 2048 10000 560 512 0.55 / 0.55
A portion of the training set is used as the background set. For global explanations, we sample a subset of the test set—either $10\%$ or $5\%$ of the full dataset—as the global reference set to estimate the expectation of the cooperative game $V(S)$. For large-scale datasets such as MIMIC-IV, a smaller subset is used for efficiency. For local explanations, a single test instance is selected. The mini-batch size for sampling $z_{i}$ is set to $10$ times the feature dimension, and the global batch size for data points is fixed at $512$.
The convergence threshold is kept consistent across all Shapley value–based methods to ensure meaningful comparisons in terms of consistency (i.e., Pearson correlation coefficient close to 1). We fix the regularization parameter at $\lambda=0.01$, and set the EMA coefficient $t$ to either $0.5$ or $0.55$, as shown in Table 1.
For baselines, we compare against SAGE algorithm proposed in [2] for global explanations, and original KernelSHAP [5] for local explanations, excluding the enhancements proposed in [16], which remain applicable to SIM-Shapley.
The MLP architectures used in our experiments are summarized in Table 2 and implemented in PyTorch. Each model contains three linear layers and two ELU activations, with varying widths across datasets. A Sigmoid activation is used for binary classification on the MIMIC-IV dataset. Both models are trained with a learning rate of 0.001 for up to 250 epochs using early stopping based on validation loss. The complete training pipeline is available in our open-source repository.
Figures 23 and 4 present additional results consistent with the main paper. Their descriptions follow the same conventions and support similar conclusions.
eTable 2: Comparison of two neural network architectures.
Layer # MLP for Airbnb MLP for MIMIC-IV
1 Linear(15, 512) Linear(56, 512)
2 ELU ELU
3 Linear(512, 512) Linear(512, 256)
4 ELU ELU
5 Linear(512, 1) Linear(256, 1)
6 Sigmoid
(a) Airbnb - Local - Consistency
(b) Airbnb - Local - Running Time
(c) Bike - Local - Consistency
(d) Bike - Local - Running Time
(e) Credit - Local - Consistency
(f) Credit - Local - Running Time
eFigure 2: Boxplot of local explanation. Comparison of Consistency (Pearson’s Correlation) and Running Time(s).
(a) Bank - Global - Consistency
(b) Bank - Global - Running Time
(c) Bike - Global - Consistency
(d) Bike - Global - Running Time
(e) Credit - Global - Consistency
(f) Credit - Global - Running Time
eFigure 3: Boxplot of global explanation. Comparison of Consistency (Pearson’s Correlation) and Running Time(s).
(a) Bank - Global Explanation
(b) Bank - Local Explanation
(c) Bike - Global Explanation
(d) Bike - Local Explanation
eFigure 4: Comparison of convergence time(s). Compare global and local interpretability methods on Bank and Bike datasets.
D.3 MIMIC-IV Experimental Details
D.3.1 Cohort Formation
The MIMIC-IV Emergency Department (MIMIC-IV-ED) dataset [37] is a publicly available resource containing over 400,000 emergency department (ED) visit episodes. Following the data extraction pipeline proposed by Xie et al. [38], we constructed a master dataset and formed our study cohort by excluding encounters with missing values in any of the 56 commonly used ED variables–such as demographics, vital signs, and comorbidities–selected for analysis.
D.3.2 Domain-Guided Interpretation of Results
To assess the clinical relevance of the feature attributions generated by SIM-Shapley and Stable-SIM-Shapley on the MIMIC-IV-ED dataset, we consulted a board-certified emergency physician. The physician independently reviewed the top-ranked features identified by our methods for predicting inpatient mortality and confirmed that they are plausible and aligned with clinical understanding. Notably, features such as vital signs, age, and comorbidity burden were consistent with known risk factors in emergency medicine. This alignment with expert knowledge supports the interpretability and practical utility of our proposed method in real-world clinical settings. These observations align with domain standards such as the Canadian Triage and Acuity Scale (CTAS) [^2] https://files.ontario.ca/moh_3/moh-manuals-prehospital-ctas-paramedic-guide-v2-0-en-2016-12-31.pdf, and comorbidity scoring systems like the Charlson Comorbidity Index [39, 40].